Trichloromethane consist of 10.089 carbon 0.83% hydrogen and 89.12% chlorine .it relatively molecular mass 119.5grams .find it molecular formula 


1
Expert's answer
2021-11-17T18:12:02-0500

Trichloromethane consist of 10.05% carbon 0.83% hydrogen and 89.12% chlorine.


Solution:

Molar mass of nitrogen (C) is 12.0 g mol-1.

Molar mass of hydrogen (H) is 1.00 g mol-1.

Molar mass of carbon (Cl) is 35.5 g mol-1.

 

Assume a 100g sample, convert the same % values to grams.

Hence,

Mass of C = w(C) × Mass of sample = 0.1005 × 100 g = 10.05 g

Mass of H = w(H) × Mass of sample = 0.0083 × 100 g = 0.83 g

Mass of Cl = w(Cl) × Mass of sample = 0.8912 × 100 g = 89.12 g

 

Convert to moles.

10.05 g C × (1 mol C / 12.0 g C) = 0.8375 mol C

0.83 g H × (1 mol H / 1.00 g H) = 0.8300 mol H

89.12 g Cl × (1 mol Cl / 35.5 g Cl) = 2.5104 mol Cl

 

Divide all moles by the smallest of the results.

C: 0.8375 / 0.8300 = 1.01 = 1

H: 0.8300 / 0.8300 = 1.00 = 1

Cl: 2.5104 / 0.8300 = 3.02 = 3


Thus, empirical formula is CHCl3


Empirical formula mass = Ar(C) + Ar(H) + 3×Ar(Cl) = 12.0 + 1.00 + 3×35.5 = 119.5

Relatively molecular mass = 119.5

n formula units = Relatively molecular mass / Empirical formula mass

n formula units = (119.5) / (119.5) = 1

Hence,

Molecular formula = Empirical formula × n formula units

Molecular formula = (CHCl3) × 1 = CHCl3


Answer: The molecular formula of trichloromethane is CHCl3